A lot is in the shape of a right triangle. The shorter leg measures 150 m. The hypotenuse is 50 m longer than the length of the longer leg. How long is the longer leg?
Answer by Loon Bajracharya(1)

Let the shorter leg be the perpendicular and the longer leg be base.
by using Pythagorean theoream,we have
h^2=p^2+b^2
now,if the hypotenious is 50m longer than longer i.e base. then h=p+50m
therefore,
(p+50m)^2=p^2+150^2
now fore the hypotenious side solve using the formula (a+b)^2
p^2+2×p×50+50^2=p^2+22500
or,100p=p^2-p^2+22500-2500
or,p=20000/100
therefore, p=longer leg=200m
SOLVED.....